OpenStudy (anonymous):

4x -4y +13z=-30 22x -22y +73z=-168 35x -35y +117z=-269

OpenStudy (anonymous):

Fun.....

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

4x -4y +13z=-30=x=13z/4+y-15/2

OpenStudy (anonymous):

Got to do elimination. Pick two pairs of equations. Eliminate z. Then, solve for x and y. Then, substitute to figure out z.

OpenStudy (anonymous):

22x -22y +73z=-168=x=-73z/22+y-84/11

OpenStudy (anonymous):

not my idea of fun, but hey, here is the answer, although this will not help you on a test http://www.wolframalpha.com/input/?i=4x+-4y+%2B13z%3D-30%2C22x+-22y+%2B73z%3D-168%2C+35x+-35y+%2B117z%3D-269

OpenStudy (anonymous):

35x -35y +117z=-269=x=-117z/35+y=269/35

hero (hero):

I would post my steps, but you would hate me afterwards.

OpenStudy (anonymous):

His solving it using matrices.

hero (hero):

No, not using matrices

OpenStudy (anonymous):

or the reduction method

hero (hero):

Using my own method

OpenStudy (anonymous):

....why don't you want to use matrices?

OpenStudy (anonymous):

Using matrices to teach this type of linear equation solving is a BAD idea. Matrix method is fine as a subsequent lesson to generalize the solution for N variables.

hero (hero):

Romero, why don't you post your matrix steps...lol

OpenStudy (anonymous):

So you rather use substitution?

OpenStudy (anonymous):

wait how is row reduction different from using matrices?

hero (hero):

I prefer using the most agonizing set of steps

OpenStudy (amistre64):

ouch, agonizing is right

hero (hero):

lol

hero (hero):

All that work just to find out no solution. I'm disappointed.

OpenStudy (amistre64):

rref a matrix would be my idea as well rref{{4, -4, +13, -30},{22, -22, +73, -168},{35, -35, +117, -269}} http://www.wolframalpha.com/input/?i=rref%7B%7B4%2C+-4%2C+%2B13%2C+-30%7D%2C%7B22%2C+-22%2C+%2B73%2C+-168%7D%2C%7B35%2C+-35%2C+%2B117%2C++-269%7D%7D infinite solutions

Directrix (directrix):

To my thinking, the solution technique may be dictated by the Instructor or the instructions. I don't know how far along CW is in his studies. If CW has never heard of matrices, then a solution based on them will mean little, I would think. Take a look at this online solver: http://www.analyzemath.com/Calculators/Calculator_syst_eq.html

hero (hero):

amistre,you should have did it by hand.